# On Uncertainty Calibration for Equivariant Functions

### v2

@@ 5.5 GENCE Invariant Lower Bound
In the case scalar case, where both the space of mean vectors $\mathcal{M}$ and the space of variance vectors $\mathcal{S}$ are one dimensional, we establish a lower bound on $\mathrm{GENCE}_{\mathrm{sq}}(h)$ using the same strategy as for ECE in Theorem˜5. The bound can be computed in a straightforward manner analogously to the classification lower bound, and the special cases for Lipschitz functions still apply. Moreover, the cases where the bound becomes vacuous are more apparent and again are analogous to what was proved for Theorem˜5.
@@
Assume that the elements of the function space $\mathscr{H}$ are invariant under the action of $G$ on $X$ and $\mathcal{M}=\mathbb{R}$, $\mathcal{S}=\mathbb{R}_{+}$. Denote by $m$ the error lower bound, $m=\underset{s\in\mathbb{R}_{+}}{\min}\int_{F_{s}}q_{s}(Gx_{s})\mathbb{V}_{Gx_{s}}[f]dx$. Then, $\mathrm{GENCE}_{\mathrm{sq}}(h)$ given by Equation˜9 is bounded below by $\int_{0}^{m}\frac{r(s)}{s^{2}}(s-m)^{2}ds$.
@@
Analogous results to Proposition˜4 and Proposition˜5 can be obtained for regression under assumption of Lipschitz continuity. Our theorem and proof strategy show that the classification and regression calibration lower bounds are both related to invariance by the minimum fiber-wise approximation error bounds when the outputs in $\mathcal{M}$ and $\mathcal{S}$ are scalar valued. Similar to ˜1, this bound is vacuous only under the condition that $f$ is constant on one each of the orbits on one of the fibers, in which case $m=0$.
@@ 6 Disentangling Aleatoric and Epistemic Uncertainty
This section details how models that are overconstrained by symmetry may misattribute the source of their uncertainties. In the previous sections, we analyzed the effects of equivariance on model miscalibration. However, our analysis did not distinguish between epistemic and aleatoric uncertainty. Understanding the breakdown between the two sources is important it allows users to determine whether they have reached the noise floor of their problem or their model is inadequate or poorly trained. While there are clear benefits to distinguishing between the two uncertainties, doing so in practice is known to be hard. Techniques such as evidential regression (as outlined in Section˜3.5) are far from perfect in their ability to disentangle these uncertainties (Ovadia et al., 2019; Valdenegro-Toro & Mori, 2022; Osband et al., 2023; Wimmer et al., 2023; Nevin et al., 2024; Jürgens et al., 2024).
With calibration error, the uncertainty predictions are often interpreted to be the epistemic uncertainty and do not consider aleatoric uncertainties. Thus, we now frame the problem of model calibration using a different metric, the aleatoric bleed, which we will now define.
@@
The aleatoric bleed is the regression error $\text{err}_{\text{reg}}(h)$ for a function space $\mathscr{H}=\{h:X\to\mathcal{S}_{\text{aleatoric}}\}$.
Put differently, the aleatoric bleed measures the epistemic mass spuriously identified as aleatoric mass. Typical ensemble spread is roughly epistemic only if the aleatoric uncertainty is correctly modeled, otherwise the uncertainties may leak both ways. Aleatoric bleed is a means of quantifying the degree to which epistemic error induced by symmetry mismatch is absorbed by the aleatoric channel of a parametric predictor. If a function $h$ estimates the aleatoric uncertainty using evidential regression as in Section˜3.5, then the aleatoric bleed is the error between predicted uncertainty vectors $\hat{s}$ with elements $\hat{s_{i}}=\beta_{i}/(\alpha_{i}-1)$ and a ground truth uncertainty $s$. We note that by Theorem˜3, the aleatoric bleed has a known lower bound.
We highlight that this bound is structural (from symmetry) rather than statistical. This is useful, because even with correct equivariance the aleatoric and epistemic separation is not identifiable without repeated-measure or instrument-noise structure. In this setting, the aleatoric uncertainty can be estimated from samples or modeled using domain knowledge, and any residual uncertainty is known to be epistemic. The only thing that can be detected is calibration failures on fibers of predicted variance, which we explored in Section˜4 and Section˜5. Lastly, we notice that the aleatoric bleed can be easily computed in cases where the ground truth aleatoric uncertainty is identically the zero vector, which we will explore in the experiments that follow.
@@ 6.1 Aleatoric Bleed Experiments
This experiment demonstrates an intuitive example where incorrect and extrinsic equivariance contributes to aleatoric bleed. We consider a model $h:\mathbb{R}^{3}\to\mathbb{R}^{3}\times\mathbb{R}^{3}$ that predicts two vector fields representing a mean and a variance prediction. That is, we predict two vectors attributed to any given point in $\mathbb{R}^{3}$ indicative of a prediction and an aleatoric uncertainty. We denote the ground truth vector field at a given point $x$ by $f(x)$, and $h$ is constrained to be $E(3)$-equivariant. Figure˜6 presents two examples of how the equivariance taxonomy can result in different levels of aleatoric bleed.
@@
For simplicity and visualization purposes, we construct our dataset with vectors in $\mathbb{R}^{3}$ with a $z$ component of 0 and we choose rotation matrices $Q$ that keep the vectors in the xyxy-plane. We provide relevant training details in Appendix˜B.
@@ 6.1.2 Chemical Properties and Aleatoric Bleed
Figure 8: Top Left: Prediction versus label for both GIN and $E(3)$-Invariant models. The two models perform similarly in terms of regression. Top Right: The difference between aleatoric and epistemic uncertainties between the two models for each label. The model’s uncertainty estimates tend to be consistent with one another. Bottom Left: A distribution of the epistemic uncertainty predictions for the two models. Bottom Right: A distribution of the aleatoric uncertainty predictions for the two models. We see that the distribution for epistemic uncertainty has a fater tail than the aleatoric uncertainty distribution.
@@ 8 Future Work
Having laid the groundwork for a first theory for equivariance and uncertainty, there are several interesting avenues for future work. An unbiased estimator for ECE and ENCE that does not depend on binning approximations would be an extremely useful contribution. Having such would enable the ability to interpret model calibration on an absolute scale instead of only being able to compare relative performances that are each dependent on discretization biasses.
@@
This work may be extended to broader experimental domains, for example, robotics and cosmology. In these areas, there is a clear intersection between uncertainty and equivariance. Uncertainty and equivariance have proved to be indespensible in robotics in particular due to cost of data collection and the safety-critical applications. In robotics, future work will examine calibration error for equivariant models for imitation learning. In particular, our approach is better suited for imitation learning that reinforcement learning. While our work can be difficult to frame in typical Q-learning setups since the optimal solution to the Bellman update function is unique, a discussion of both equivariance and uncertainty quantification lends itself naturally to behavior cloning tasks (Florence et al., 2022).
Another closely related experimental domain is cosmological large-scale structure. In particular, future work may assess calibration error with symmetry-preserving models using the benchmark released in Balla et al. (2024). The appeal of this benchmark is that it includes graph-level predictions on $\Lambda CDM$ (Ryden, 2016; Carroll, 2019) cosmological parameters $\Omega_{m}$ and $\sigma_{8}$. Moreover, the $\Omega_{m}$ and $\sigma_{8}$ parameters are extremely important measurements detailing the matter density of our universe. In cosmology, these predictions are more commonly phrased as constraints on posterior distributions rather than point estimates (e.g., Collaboration: et al., 2016; Abbott et al., 2022). Therefore, it is reasonable to apply our framework to this dataset and in particular assess if equivariance can help models distinguish between epistemic and aleatoric uncertainties.
@@ 9 Discussion and Conclusions
The core conclusions are best explained through the lens of model mispecification, and highlight how equivariant neural networks can fail to meet their calibration objectives on datasets that do not share the same symmetries. We highlight how when equivariant model assumptions are violated, i.e. cases of incorrect and extrinsic equivariance where the model is either over-constrained by symmetry or forced to treat out-of-distribution points as in-distribution, model calibration for both classification and regression tasks is provably worse. Our experiments support these conclusions as well. In the cases of the swiss roll and vector field regression experiments, both incorrect and extrinsic equivariance not only make the model less accurate, but also poorly calibrated. In the case of the vector field regression experiment we also saw that incorrect and extrinsic equivariance contributed to aleatoric bleed. By contrast, we have shown that a model with correct equivariance is not necessarily better calibrated than a similarly sized non-equivariant baseline. As illustrated by the galaxy morphology classification and scalar-valued chemical property prediction experiments, equivariance is not strong enough to help a model become well calibrated nor is it strong enough to prevent aleatoric bleeding. It was only for the highly synthetic vector regression experiment on the sinusoidal dataset that the introduction of correct equivariance was able to significantly improve the raw performance and prevent aleatoric bleeding. The vector regression result is especially interesting in light of the galaxy morphology classification experiment, which showed that correct equivariance can help a model perform better without necessarily making it better calibrated.
@@ 10 Reproducability Statement
Our codebase containing all of our experiments, as well as the instructions to reproduce our results, is publicly available at https://github.com/EdwardBerman/EquivariantUQBounds.
@@ Acknowledgments
The authors thank Shubhendu Trivedi for thoughtful discourse and detailed review on an initial draft of the manuscript. E.B. and J.G. acknowledge feedback and overwhelming support from students in Northeastern University’s Mathematics course. E.B. additionally thanks the Biomarkers team at AstroAI (Garraffo, 2024) for helpful discussions. Finally, E.B. and J.G. thank members of the Geometric Learning Lab for helpful discussions and welcoming support.
@@ Appendix A Iterated Integration
If we assume that $\cup_{g_{1}\neq g_{2}}(g_{1}F\cap g_{2}F)$ has measure 0 and $F$ and $Gx$ are differentiable manifolds, then we may lift an integral $Gx$ to itself. Denote the identification of the orbit $Gx$ and coset space $G/G_{x}$ with respect to the stabilizer $G_{x}=\{g:gx=x\}$ by $a_{x}:G/G_{x}\to Gx$. Then we have
@@ Appendix B Vector Regression Setup
Our training of the $E(3)$-equivariant neural network uses e3nn_jax (Geiger et al., 2022; Geiger & Smidt, 2022; Kondor et al., 2018; Weiler et al., 2018; Thomas et al., 2018). The MLP baseline is built entirely with Flax (Heek et al., 2024). The models are trained for a minimum of $10$ epochs for a maximum number of $100$, with early stopping if the validation loss stops improving after $5$ epochs. We train on $2000$ generated samples. We train using both a $\beta$-NLL loss (Seitzer et al., 2022) and an MSE loss, equally weighted, with $\beta=1$. The $\beta$-NLL loss is given by

### v4

@@ 5.4 GENCE Invariant Lower Bound
In the scalar case, where both the space of mean vectors $\mathcal{M}$ and the space of variance vectors $\mathcal{S}$ are one dimensional, we establish a lower bound on $\mathrm{GENCE}_{\mathrm{sq}}(h)$ using the same strategy as for ECE in theorem˜5. The bound can be computed in a straightforward manner analogously to the classification lower bound, and the special cases for Lipschitz functions still apply. Moreover, the cases where the bound becomes vacuous are more apparent and again are analogous to what was proved for theorem˜5.
@@
Assume that the elements of the function space $\mathscr{H}$ are invariant under the action of $G$ on $X$ and $\mathcal{M}=\mathbb{R}$, $\mathcal{S}=\mathbb{R}_{+}$. Denote by $m$ the error lower bound, $m=\underset{s\in\mathbb{R}_{+}}{\min}\int_{F_{s}}q_{s}(Gx_{s})\mathbb{V}_{Gx_{s}}[f]dx$. Then, $\mathrm{GENCE}_{\mathrm{sq}}(h)$ given by equation˜10 is bounded below by $\int_{0}^{m}\frac{r(s)}{s^{2}}(s-m)^{2}ds$.
@@
Analogous results to proposition˜4 and proposition˜5 can be obtained for regression under assumption of Lipschitz continuity. Our theorem and proof strategy show that the classification and regression calibration lower bounds are both related to invariance by the minimum fiber-wise approximation error bounds when the outputs in $\mathcal{M}$ and $\mathcal{S}$ are scalar valued. Similar to ˜1, this bound is vacuous only under the condition that $f$ is constant on each of the orbits on one of the fibers, in which case $m=0$.
5.5 Example of GENCE Upper Bound for Invariant Models
Example 7.
The goal of this example is to show how the GENCE bound can be computed in a small synthetic example and to illustrate how the bound can be tighter or looser depending on the norm of the variances. In this example, we consider a set $X$ of five point clouds. We calculate the GENCE upper bound assuming a minimized regression error. We assume that the function space $\mathscr{H}$ has $E(2)$-invariance. As in section˜4.7, this dataset is adversarial to the entire function space $\mathscr{H}$ given that the ground truth function $f$ is not $E(2)$-invariant, as indicated in figure˜5. The dataset contains some duplicates of the point clouds up to transformations in $E(2)$. Different point cloud orientations are notated with ++ and $\times$.
We now compute the upper bound by computing the regression error lower bound along each fiber. We start with the first fiber (rows 1 and 2 in figure˜5). Since $h$ is $E(2)$-invariant, $h$ is able to fully fit the function $f$ on the first orbit (row 1) containing two rotated versions of $a$. Thus, the regression error is zero. For the second row, $b$ and $c$ are equally probable, and the output of $h$ that minimizes the regression error is just the average of the two lines. To compute the error on the second orbit (row 2), we also need to compute the integrated density:
$q(Gx_{s})$ $=\frac{p(b)+p(c)}{p(a(+))+p(a(\times))+p(b)+p(c)}=0.5.$
Applying equation˜12 gives $\text{err}_{\text{reg}}(h,s_{1})=0.5\cdot\mathbb{V}_{Gx_{s_{1}}}[f]=\pi/8$. Now, as with the first orbit, the minimizing regression error on the third orbit (row 3) is zero, since there is only one element in $X$ we can fit. So, $\text{err}_{\text{reg}}(h,s_{2})=0$. Finally, that the probability of sampling an element on the fiber of $s_{1}$ is the same as the probability of sampling an element on $s_{2}$. Thus, by theorem˜2,
[…]
The upper bound for GENCE is thus $1+\frac{\pi/8}{\lVert\sqrt{\frac{2}{\pi}s_{1}}\rVert_{2}^{2}}$. We can see that in the limit as $\lVert\sqrt{\frac{2}{\pi}s_{1}}\rVert_{2}^{2}$ goes to infinity, the upper bound on GENCE becomes $1$. Alternatively, in the limit as $\lVert\sqrt{\frac{2}{\pi}s_{1}}\rVert_{2}^{2}$ goes to 0, the upper bound on GENCE diverges. The interpretation of the latter is that if the model is extremely confident, then any deviations from the mean prediction represents similarly extreme miscalibration.
@@ 6 Disentangling Aleatoric and Epistemic Uncertainty
This section details how models that are overconstrained by symmetry may misattribute the source of their uncertainties. In the previous sections, we analyzed the effects of equivariance on model miscalibration. However, our analysis did not distinguish between epistemic and aleatoric uncertainty. Understanding the breakdown between the two sources is important because it allows users to determine whether they have reached the noise floor of their problem or their model is inadequate or poorly trained. While there are clear benefits to distinguishing between the two uncertainties, doing so in practice is known to be hard. Techniques such as evidential regression (as outlined in section˜3.5) are far from perfect in their ability to disentangle these uncertainties (Ovadia et al., 2019; Valdenegro-Toro and Mori, 2022; Osband et al., 2023; Wimmer et al., 2023; Nevin et al., 2024; Jürgens et al., 2024).
6.1 Preliminaries
With calibration error, the uncertainty predictions are often interpreted to be the epistemic uncertainty and do not consider aleatoric uncertainties. Thus, we now frame the problem of model calibration using a different metric, the aleatoric bleed, which we will define after detailing the relevant background from section˜3.5.
Recall the definitions of aleatoric and epistemic uncertainty in section˜3.5, which we repeat here for convenience.
Aleatoric Uncertainty:
Aleatoric uncertainty refers to the irreducible part of the uncertainty. Given spaces $X$ and $Y$ and an instance $x_{q}\in X$, the aleatoric uncertainty is the spread in $p(y|x_{q})$.
Epistemic Uncertainty:
Model uncertainty and approximation uncertainty, on the other hand, are subsumed under the notion of epistemic uncertainty. Let the spaces $X$ and $Y$ be the same as before. Let $l:Y\times Y\to\mathbb{R}$ be the loss function and let $f^{*}$ be the associated point-wise Bayes predictor defined as
$f^{*}(x):=\underset{\hat{y}\in Y}{\arg\min}\int_{Y}l(y,\hat{y})dP(y|x).$ (14)
Epistemic uncertainty is the uncertainty due to the lack of knowledge of the perfect predictor equation˜14.
@@ 6.1 Preliminaries
The aleatoric bleed is the regression error $\text{err}_{\text{reg}}(h)$ for a function space $\mathscr{H}=\{h\colon X\to\mathcal{S}_{\text{aleatoric}}\}$.
Put differently, the aleatoric bleed measures the epistemic mass spuriously identified as aleatoric mass. Typical ensemble spread is roughly epistemic only if the aleatoric uncertainty is correctly modeled, otherwise the uncertainties may leak both ways. Aleatoric bleed is a means of quantifying the degree to which epistemic error induced by symmetry mismatch is absorbed by the aleatoric channel of a parametric predictor.
Again for convenience, we now recap the specific parameterization of the uncertainties via evidential regression that was outlined in section˜3.5. Given data points $\left(y_{1},\ldots,y_{n}\right)\sim\mathcal{N}(\mu,\sigma^{2})$, we may impose priors $\mu\sim\mathcal{N}(\gamma,\sigma^{2}\nu^{-1})$, $\sigma^{2}\sim\Gamma^{-1}(\alpha,\beta)$ where $\Gamma(\cdot)$ is the gamma function, $m=(\gamma,\nu,\alpha,\beta)$, and $\gamma\in\mathbb{R}$, $\nu>0$, $\alpha>1$, $\beta>0$. One can then show that $p(y_{i}|m)=St(y_{i};\gamma,\beta(1+\nu)/\alpha\nu,2\alpha)$, where the St distribution has probability density given by
Parameterizing the Student’s t distribution as a four parameter family is useful because it lends itself to the following definitions of the prediction, aleatoric uncertainty, and epistemic uncertainty:
$\mathbb{E}[\mu]$ $=$ $\gamma\quad\text{(Prediction)}$
If a function $h$ estimates the aleatoric uncertainty using evidential regression, then the aleatoric bleed is the error between predicted uncertainty vectors $\hat{s}$ with elements $\hat{s_{i}}=\beta_{i}/(\alpha_{i}-1)$ and a ground truth uncertainty $s$. We note that by theorem˜3, the aleatoric bleed has a known lower bound. Our choice of evidential regression as opposed to other uncertainty decompositions such as mutual information is motivated by the issues laid out in Wimmer et al. (2023). While evidential regression can struggle to perform uncertainty decomposition due to convergence issues (Bengs et al., 2022; Meinert et al., 2023; Jürgens et al., 2024), we expect it to perform better in the cases without symmetry mismatch and with a known ground truth aleatoric uncertainty and ultimately provide a strong baseline performance.
We highlight that this bound is structural (from symmetry) rather than statistical. This is useful because even with correct equivariance, the aleatoric and epistemic separation is not identifiable without repeated-measure or instrument-noise structure. In this setting, the aleatoric uncertainty can be estimated from samples or modeled using domain knowledge, and any residual uncertainty is known to be epistemic. The only thing that can be detected is calibration failures on fibers of predicted variance, which we explored in section˜4 and section˜5. Lastly, we notice that the aleatoric bleed can be easily computed in cases where the ground truth aleatoric uncertainty is identically the zero vector, which we will explore in the experiments that follow.
@@ 6.2 Aleatoric Bleed Experiments
This experiment demonstrates an intuitive example where incorrect and extrinsic equivariance contributes to aleatoric bleed. Consider a model $h:\mathbb{R}^{3}\to\mathbb{R}^{3}\times\mathbb{R}^{3}$ that predicts two vector fields representing a mean and a variance prediction. That is, we predict two vectors attributed to any given point in $\mathbb{R}^{3}$ indicative of a prediction and an aleatoric uncertainty. We denote the ground truth vector field at a given point $x$ by $f(x)$, and $h$ is constrained to be $E(3)$-equivariant. figure˜6 presents two examples of how the equivariance taxonomy can result in different levels of aleatoric bleed.
@@
For simplicity and visualization purposes, our dataset consists of vectors in $\mathbb{R}^{3}$ with a $z$ component of 0, and we choose rotation matrices $Q$ that keep the vectors in the $xy$-plane. We provide relevant training details in appendix˜B.
@@ 6.2.2 Chemical Properties and Aleatoric Bleed
Figure 8: Top Left: Prediction versus label for both GIN and $E(3)$-Invariant models. The two models perform similarly in terms of regression. Top Right: The difference between aleatoric and epistemic uncertainties between the two models for each label. The model’s uncertainty estimates tend to be consistent with one another. Bottom Left: A distribution of the epistemic uncertainty predictions for the two models. Bottom Right: A distribution of the aleatoric uncertainty predictions for the two models. We see that the distribution for epistemic uncertainty has a fatter tail than the aleatoric uncertainty distribution.
@@ 8 Future Work
Having laid the groundwork for a first theory for equivariance and uncertainty, there are several interesting avenues for future work. An unbiased estimator for ECE and ENCE that does not depend on binning approximations would be an extremely useful contribution. Having such would enable the ability to interpret model calibration on an absolute scale instead of only being able to compare relative performances that are each dependent on discretization biases.
@@
This work may be extended to broader experimental domains, for example, robotics and cosmology. In these areas, there is a clear intersection between uncertainty and equivariance. Uncertainty and equivariance have proved to be indispensable in robotics in particular due to cost of data collection and the safety-critical applications. In robotics, future work will examine calibration error for equivariant models for imitation learning. In particular, our approach is better suited for imitation learning than reinforcement learning. While our work can be difficult to frame in typical Q-learning setups since the optimal solution to the Bellman update function is unique, a discussion of both equivariance and uncertainty quantification lends itself naturally to behavior cloning tasks (Florence et al., 2022).
[…]
@@ 9 Discussion and Conclusions
The core conclusions are best explained through the lens of model mispecification, and highlight how equivariant neural networks can fail to meet their calibration objectives on datasets that do not share the same symmetries. We highlight how when equivariant model assumptions are violated, i.e., cases of incorrect and extrinsic equivariance where the model is either over-constrained by symmetry or forced to treat out-of-distribution points as in-distribution, model calibration for both classification and regression tasks is provably worse. Our experiments support these conclusions as well. In the cases of the swiss roll and vector field regression experiments, both incorrect and extrinsic equivariance not only make the model less accurate, but also poorly calibrated. In the case of the vector field regression experiment we also saw that incorrect and extrinsic equivariance contributed to aleatoric bleed. By contrast, we have shown that a model with correct equivariance is not necessarily better calibrated than a similarly sized non-equivariant baseline. As illustrated by the galaxy morphology classification and scalar-valued chemical property prediction experiments, equivariance is not strong enough to help a model become well calibrated nor is it strong enough to prevent aleatoric bleeding. It was only for the highly synthetic vector regression experiment on the sinusoidal dataset that the introduction of correct equivariance was able to significantly improve the raw performance and prevent aleatoric bleeding. The vector regression result is especially interesting in light of the galaxy morphology classification experiment, which showed that correct equivariance can help a model perform better without necessarily making it better calibrated.
@@ 10 Reproducability Statement
Our codebase containing all of our experiments, as well as the instructions to reproduce our results, is publicly available at github.com/EdwardBerman/EquiUQ.
@@ Acknowledgments
The authors thank Shubhendu Trivedi for thoughtful discourse and detailed review on an initial draft of our manuscript. E.B. and J.G. acknowledge the feedback and overwhelming support from students in Northeastern University’s Mathematics Research Capstone course. E.B. additionally thanks the Biomarkers team at AstroAI (Garraffo, 2024). Finally, E.B., J.G., and M.P. thank the members of the Geometric Learning Lab for helpful discussions and welcoming support.
R.W. would like to acknowledge support from NSF Grants $2442658$ and $2134178$. This work is supported by the National Science Foundation under Cooperative Agreement PHY-$2019786$ (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, http://iaifi.org/).
@@ Appendix A Iterated Integration
If we assume that $\cup_{g_{1}\neq g_{2}}(g_{1}F\cap g_{2}F)$ has measure 0 and $F$ and $Gx$ are differentiable manifolds, then we may lift an integral $Gx$ to itself. Denote the identification of the orbit $Gx$ and coset space $G/G_{x}$ with respect to the stabilizer $G_{x}=\{g\colon gx=x\}$ by $a_{x}\colon G/G_{x}\to Gx$. Then we have
@@ Appendix B Vector Regression Setup
Our training of the $E(3)$-equivariant neural network uses e3nn_jax (Geiger et al., 2022; Geiger and Smidt, 2022; Kondor et al., 2018; Weiler et al., 2018; Thomas et al., 2018). The MLP baseline is built entirely with Flax (Heek et al., 2024). The models are trained for a minimum of $10$ epochs for a maximum number of $100$, with early stopping if the validation loss stops improving after $5$ epochs. We train on $2000$ generated samples. We train using both a $\beta$-NLL loss (Seitzer et al., 2022) and an MSE loss, equally weighted, with $\beta=1$. The $\beta$-NLL loss is given by